OpenStudy (anonymous):

Help! I will grant a medal! Can someone explain how to solve this? y-5/3 = 1

OpenStudy (anonymous):

I think the answer is 8, but I don't know the steps to solve it ( I used mental math to guess 8)

OpenStudy (anonymous):

@KirbyLegs @Wariety

OpenStudy (anonymous):

I would agree :) I love answer your problems! You get them correct a lot too . . .

OpenStudy (anonymous):

is the equation \[\frac{y-5}{3}=1 or y-\frac{5}{3}=1\]

OpenStudy (anonymous):

\[\frac{ y - 5 }{ 3 } = 1\] @Wariety

OpenStudy (anonymous):

Then u are correct.

OpenStudy (anonymous):

I know, I did mental math to solve that. I just don't know the steps to solve it.

OpenStudy (anonymous):

^that is the problem

OpenStudy (anonymous):

The steps would be to first multiply each side by 3 to get rid of that ugly division on the left. That would leave you with y-5 = 3

OpenStudy (anonymous):

Then add five to both sides to do away the -5 on the left. That will leave you with y=8

OpenStudy (anonymous):



OpenStudy (anonymous):

@Wariety Thanks

OpenStudy (anonymous):

Call me paranoid, but in terms of presentation, I would not show "3+5=8", cause that doesn't really prove that y=8. I would have shown it like this: \[\frac{y-5}{3}=1\] \[\frac{y-5}{3}\times3=1\times3\] \[y-5(+5)=3(+5)\] \[y=3+5\] \[y=8\]

OpenStudy (anonymous):

I agree. I just thought I would include it anyways

OpenStudy (anonymous):

Step 3 isnt really necessary